# Object-Centric Image to Video Generation with Language Guidance

### v1

@@ Appendix E Additional Results
Figure 9: Comparison between SAVi and our Extended DINOSAUR (Ext. DINO) decomposition modules in reconstructing an CLIPort video sequence.
E.1 SAVi vs. DINO
Current object-centric approaches for video generation are limited to relatively simple synthetic datasets. We attribute this limitation primarily to the object-centric modules used for learning object representations. Motivated by this observation, we extend DINOSAUR (Seitzer et al., 2023) to handle video data and reconstruct images effectively.
To demonstrate the significance of the object-centric module in scaling to more complex datasets, we input the same video sequence to both SAVi and our extended DINOSAUR trained on CLIPort. As illustrated in Fig. 9, SAVi struggles to accurately model objects on the 2D plane, whereas our proposed object-centric model successfully reconstructs the scene, closely resembling the input. The visual features extracted by the DINOv2 (Oquab et al., 2023) encoder contain high-level semantic information, and during training, the slots are specifically optimized to efficiently encode this information. This design enables the model to scale and handle more complex object-centric video data effectively.
E.2 Robustness to Number of Objects
We evaluate the performance of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set consisting of scenes with a larger number of objects than encountered during training, i.e. the 2D table plane contains 8 instead of 6 objects. This evaluation demonstrates the robustness of models when generalizing to scenes with more objects. The quantitative results are presented in Tab. 5. We report both the video generation performance and the performance drop relative to the original evaluation set.
We observe that our model outperforms $\text{MAGE}_{\text{DINO}}$ on every metric, for both prediction horizons. Notably, TextOCVP demonstrates significantly higher robustness to scenes with more objects, as reflected by the smaller decline in performance for the perceptual LPIPS metric. This result is further illustrated in Fig. 10, where we show a qualitative comparison of video generations for scenes with 8 objects. As observed, our model correctly illustrates the motion described in the text instructions, while $\text{MAGE}_{\text{DINO}}$ fails to generate accurate sequences, missing the target block, a behaviour observed in various examples.
These results highlight once again the effectiveness of object-centric representations in image-to-video generation, as TextOCVP is able to generalize to scenes with more objects by simply increasing the number of slots.
@@ E.2 Robustness to Number of Objects
$\text{MAGE}_{\text{DINO}}$ 0.929 (-1.2%) 0.088 (-37.5%) 0.920 (-1.2%) 0.094 (-25.3%)
TextOCVP 0.936 (-1.5%) 0.076 (-22.6%) 0.921 (-1.1%) 0.090 (-15.4%)
@@
Figure 10: Qualitative evaluation of $\text{MAGE}_{\text{DINO}}$ and TextOCVP on a CLIPort sequence with more objects than those seen during training. TextOCVP correctly generates a sequence where the robot picks up the gray block and places it in the brown bowl, while $\text{MAGE}_{\text{DINO}}$ fails by missing the target bowl.
E.3 Text-to-Slot Attention Visualizations
An additional advantage of using object-centric representations is the improved interpretability. This can be shown in the visualizations of text-to slot attention weights, which help us understand how the textual information influences and guides the model predictions. First, in Fig. 11, we visualize the text-to-slot attention weights for different cross-attention heads for a single slot that represents the rotating red cube. We observe that the slot attends to relevant text tokens from the input, such as the object shape, size and the action taking place.
@@ E.3 Text-to-Slot Attention Visualizations
Figure 11: Visualization of the text-to-slot attention between an object and the textual caption. We visualize the text-to-slot attention weights for different heads. Our predictor module focuses on relevant text tokens to predict the scene dynamics, including the object size, shape or the action.
In Fig. 12, we additionally visualize the text-to-slot attention weights, averaged across attention heads, for different slots in a CATER sequence. We observe that slots that represent objects in the textual description attend to relevant text tokens, such as their target coordinate locations. These results demonstrate that the text-to-slot attention mechanism effectively aligns the textual information with the object-centric representations, enabling the model to generate accurate video sequences based on the given text instructions, while providing superior interpretability.
[…]
‘the cone is sliding to (-3, -3).
the snitch is picked up and placed to (-1, 2 ).’
Refer to captionRefer to caption1.00.50.0Refer to caption
Slot 3
the
cone
is
sliding
to
-3
-3
the
snitch
is
picked
up
and
placed
to
-1
2
Refer to captionRefer to caption1.00.50.0Refer to caption
Slot 4
the
cone
is
sliding
to
-3
-3
the
snitch
is
picked
up
and
placed
to
-1
2
Refer to captionRefer to caption1.00.50.0Refer to caption
Slot 6
the
cone
is
sliding
to
-3
-3
the
snitch
is
picked
up
and
placed
to
-1
2
Figure 12: Text-to-slot attention weights, averaged across attention heads, for different objects in the sequence. Slots that represent objects in the textual description attend to relevant text tokens, such as their target locations.
E.4 Qualitative Evaluations
Figs. 13 and 14  show qualitative evaluations on CATER in which both MAGE and TextOCVP successfully generate a sequence following the instructions from the textual description. Fig. 15 illustrates an example where MAGE fails to generate a correct sequence, while TextOCVP successfully completes the task described by the text.
Figs. 16 and 17  show examples of TextOCVP’s control over the predictions. In both sequences, TextOCVP generates a correct sequence given the text instructions, and seamlessly adapts its generations to a modified version of the textual instructions.

### v2

@@ Appendix F Additional Results
Figure 8: Comparison between SAVi and our Extended DINOSAUR (Ext. DINO) decomposition modules for reconstructing a CLIPort sequence. We visualize the reconstructed frames, as well as the slot masks obtained by Extended DINOSAUR. SAVi fails to reconstruct most objects, whereas Extended DINOSAUR accurately reconstructs the scene, while representing each object.
F.2 Quantitative Comparison
In the main paper, we present quantitative evaluations on the CATER and CLIPort datasets using three evaluation metrics, shown in two separate tables.
For completeness, we report in Table 6 a text-guided video prediction evaluation of TextOCVP and multiple baseline models on CATER and CLIPort using four distinct evaluation metrics for different prediction horizons.
[…]
F.3 Computational Efficiency
In Table 7 we report the average inference time for $T=1$ and $T=9$ frame predictions on CATER and CLIPort using a NVIDIA-A6000-48Gb GPU.
[…]
Table 7: Average inference (Inf.) time for $T=1$ and $T=9$ frame predictions on CATER and CLIPort.
CATER Inf. [ms] CLIPort Inf. [ms]
Model $T=1$ $T=9$ $T=1$ $T=9$
MAGE 15.4 $\pm$ 1 111.1 $\pm$ 11 34.4 $\pm$ 3 198.8 $\pm$ 2
[…]
@@ F.3 Computational Efficiency
$\text{MAGE}_{\text{DINO}}$ 0.929 -1.2% 0.088 -37% 0.920 -1.2% 0.094 -25%
TextOCVP 0.936 -1.5% 0.076 -23% 0.921 -1.1% 0.090 -15%
F.4 Robustness to Number of Objects
We evaluate the performance of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set consisting of scenes with a larger number of objects than encountered during training, i.e. the 2D table plane contains 8 objects instead of the default 6 objects. This evaluation benchmarks the robustness of models when generalizing to scenes with a previously unseen number of objects. The quantitative results are presented in Table 8. We report both the video prediction performance, as well as the performance drop relative to the original evaluation set.
We observe that our model outperforms $\text{MAGE}_{\text{DINO}}$ on every metric for both prediction horizons. Notably, TextOCVP demonstrates significantly higher robustness to scenes with more objects, as reflected by the smaller decline in performance for the perceptual LPIPS metric. This result is further illustrated in Fig. 9, where we show qualitative comparisons of video generations for scenes with 8 objects. As observed, our model correctly predicts sequences following the motion described in the text instructions, whereas $\text{MAGE}_{\text{DINO}}$ fails to generate accurate sequences according to the descriptions.
These results further demonstrate the effectiveness of object-centric representations for video prediction, as TextOCVP is able to generalize to scenes with more objects by simply increasing the number of slots. This flexibility is enabled by initializing object slots via sampling from a learned Gaussian, allowing the use of a variable number of slots at test time while breaking symmetry and preserving permutation invariance. Although the object slots do not specialize during training, they reliably bind to meaningful entities through iterative attention during inference. This results in robust and scalable scene decomposition, enabling accurate modeling of complex scenes with varying object counts.
@@ F.4 Robustness to Number of Objects
t=1t=1551010202030304040a)
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the yellow block in the gray bowl Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
$\text{MAGE}_{\text{DINO}}$
TextOCVP
t=1t=110102020303040405050b)
Figure 9: Qualitative evaluation of $\text{MAGE}_{\text{DINO}}$ and TextOCVP on CLIPort sequences with more objects than those seen during training. TextOCVP correctly generates sequences according to the text instructions, whereas $\text{MAGE}_{\text{DINO}}$ misses the target bowl.
Table 9: Impact of visual artifacts on TextOCVP’s performance on CLIPort. Cropping the artifact-prone bottom part of the predicted frames leads to a notable improvement in TextOCVP’s performance.
CLIPort1→9 CLIPort1→19
@@
Full Image 0.950 0.062 0.931 0.078
Excluded Bottom 0.953 0.050 0.932 0.069
F.5 Impact of Visual Artifacts
As already discussed in the main paper, TextOCVP occasionally generates visual artifacts on the CLIPort dataset, most noticeably in the bottom-center region of the frame, where blurry patches often appear.
To assess the impact of these visual artifacts on TextOCVP’s quantitative performance, we evaluate our model after removing the last bottom rows from the predicted frames—an area that contains only background pixels and the artifacts. As shown in Table 9, removing these rows leads to improved results, particularly for the perceptual LPIPS metric. These findings verify that TextOCVP generates future frame predictions that closely follow the text description, and that its overall performance is underestimated due to localized visual artifacts. Mitigating such artifacts through improved architectures or training strategies presents an interesting direction for future work.
F.6 Text-To-Slot Attention
An additional advantage of using object-centric representations is improved interpretability. This is illustrated by visualizing the text-to-slot attention weights, which reveal how textual information guides the model’s predictions.
In Fig. 11, we visualize the text-to-slot attention weights from different cross-attention heads for a single slot that represents a rotating red cube. We observe that the slot attends to distinct relevant text tokens from the language description on each attention head, including the object shape (head 2), size (head 1) and the action taking place (head 4).
‘put the gray block in the brown bowl
[…]
TextOCVPt=1t=12020303040405050
Figure 10: Qualitative result on CLIPort. Top row shows ground truth frames. TextOCVP completes the pick-and-place task, whereas MAGE fails to predict the robot motion.
F.7 Additional Qualitative Evaluations
Fig. 10 shows an example where we evaluate TextOCVP and $\text{MAGE}_{\text{DINO}}$ for text-guided video prediction over a long prediction horizon of 50 frames. $\text{MAGE}_{\text{DINO}}$ fails to complete the task outlined in the textual description, as it stops generating consistent robot motion after 30 frames. In contrast, TextOCVP successfully predicts future frames where the robot completes the pick-and-place task.
Figs. 12 and 13  show qualitative evaluations on CATER in which both MAGE and TextOCVP successfully predict sequence continuations following the instructions from the textual description.
Fig. 14 illustrates an example where MAGE fails to generate a correct sequence, while TextOCVP successfully completes the task described by the text.
Figs. 15 and 16  show examples of TextOCVP’s control over the predictions. In both sequences, TextOCVP generates a correct sequence given the text instructions, and seamlessly adapts its generations to a modified version of the textual instructions.
@@ F.7 Additional Qualitative Evaluations
Figure 11: Visualization of the text-to-slot attention weights between an object slot and the textual tokens. We visualize the text-to-slot attention weights for four different heads. Our predictor module focuses on distinct and relevant text tokens on each head, including the object size, shape or the action.